# No-Regret Learning of Nash Equilibrium for Black-Box Games via Gaussian Processes

### v1

@@ 1 Introduction
Given the lack of agents’ utility information and the expensive query mentioned above, this paper studies efficient no-regret learning of the NE for black-box games via GP. To the best of our knowledge, there were no existing GP algorithms for learning NE with a known no-regret guarantee. The key innovation in our work is the design of a novel GP objective specifically for NE learning. Specifically, we characterize the equilibrium computation as an optimization problem involving an unknown loss function. This function represents the maximum utility gain that agents can achieve by deviating from the given strategy. Notably, reaching a zero value of this function corresponds to the NE, corresponding to a scenario where no agent can improve their utility by changing their strategy given the strategies of the others.
@@ 2.1 Bayesian optimization approach
The most closely related line of research focuses on addressing game-theoretic models that are computationally expensive to evaluate using Bayesian Optimization (BO) techniques. Al-Dujaili et al. [2018] proposed a method to find equilibria for such games in a sequential decision-making framework using BO. Specifically, they introduced the game-theoretical regret of a strategy profile $\bm{x}$ as the most utility any agent $i$ can gain by deviating from $x_{i}$ to any strategy in $\mathcal{X}_{i}$. The authors employ BO in minimizing an approximation of the game-theoretic regret to approximate the pure strategy NE. The performance in terms of game-theoretical regret of the proposed method is validated on a collection of synthetic games by comparison with some recent algorithms.
@@
Marchesi et al. [2020] proposed a multi-arm bandit algorithm on top of the Gaussian processes and offers theoretical justification. Our work differentiates from two perspectives. First, Marchesi et al. [2020] focused on two-player zero-sum games while our work allows multi-player normal-form games. Second, the regret analysis in Marchesi et al. [2020] relied on a suboptimal gap in the denominators of the regret bound. As discussed by Lattimore and Szepesvári [2020], the major problem with this dependency is that this gap, in practice, could be arbitrarily small and downgrades the practicality of the resulting regret analysis. at the same time, our theoretical results of the regret bound rely on the maximum mutual information of GP instead and is gap-independent.
@@ 2.2 Other online learning algorithms
Similar to the work by Aprem and Roberts [2021], Chapman et al. [2013] also studied convergence to Nash equilibria in potential games with rewards that are initially unknown. Different from the Bayesian optimization approach, they proposed a multi–agent version of Q-learning to estimate the reward functions using novel forms of the $\epsilon$–greedy learning policy. Jordan [1991] studied Bayesian learning of equilibrium, assuming each agent knows their utility information but not others. This work is also related to learning other equilibrium concepts in game theory and Bayesian optimization with multiple structured utility functions, we refer to Appendix A for more detailed discussions and comparisons.
@@ 4.3 Efficient High-dimensional Optimization through ROI Reduction
when setting ${\hat{\mathcal{X}}}^{0}=\mathcal{X}$. This means that the ROI identification is actually a hierarchical filtering of the search space and is accelerated by its continuing shrinkage.
There is no guarantee of the ROI shrinkage rate, potentially making its performance unstable in High-Dimensional BO (HDBO) tasks. There are several potential solutions. There are chances to incorporate existing orthogonal HDBO techniques, including sparse GP [McIntire et al., 2016, Moss et al., 2023] and dimension reduction for BO [Song et al., 2022, Wang et al., 2016, Letham et al., 2020, Munteanu et al., 2019, Papenmeier et al., 2022]. However, the methods require additional structural assumptions that do not necessarily hold in NE discovery and, therefore, require cautiousness depending on the application.
@@ A.1 Learning Stackelberg games
More broadly, our research is also related to learning the equilibrium in the game theory. Besides the Nash equilibrium studied in this paper, another well-studied game is the Stackelberg game [Von Stengel and Zamir, 2004, Gan et al., 2023]. Specifically, Stackelberg games model a two-step sequential decision-making process between two agents, a leader and a follower. This canonical model for strategic leader-follower interactions has been adopted for many applications in the real world, such as contract design, optimal pricing, security resource allocation, and optimal traffic routing [Bolton and Dewatripont, 2004, Dawkins et al., 2021, Paruchuri et al., 2008, Roth et al., 2016, Roughgarden, 2001, Yang et al., 2014]. Learning the Stackelberg equilibrium has also been extensively studied in the literature [Letchford et al., 2009, Peng et al., 2019, Dawkins et al., 2022, Han et al., 2023], it would be interesting to study the learning of Stackelberg game equilibria via Gaussian Processes.

### v2

@@ 1 Introduction
Given the lack of agents’ utility information and the expensive query mentioned above, this paper studies efficient no-regret learning of the NE for black-box games via GP. To the best of our knowledge, there were no existing GP algorithms for learning NE with a known no-regret guarantee. The key innovation in our work is the design of a novel GP objective specifically for NE learning. Specifically, we characterize the equilibrium computation as an optimization problem involving an unknown loss function. This function represents the maximum utility gain that agents can achieve by deviating from the given strategy. Notably, reaching a zero value of this function corresponds to the NE, a scenario where no agent can improve their utility by changing their strategy given the strategies of others.
@@ 2.1 Bayesian optimization approach
The most closely related line of research focuses on addressing game-theoretic models that are computationally expensive to evaluate using Bayesian Optimization (BO) techniques. Al-Dujaili et al. [2018] proposed a method to find equilibria for such games in a sequential decision-making framework using BO. Specifically, they introduced the game-theoretical regret of a strategy profile $\bm{x}$ as the most utility any agent $i$ can gain by deviating from $x_{i}$ to any strategy in $\mathcal{X}_{i}$. The authors employ BO to minimize an approximation of the game-theoretic regret and approximate the pure strategy NE. The performance in terms of game-theoretical regret of the proposed method is validated on a collection of synthetic games by comparison with some recent algorithms.
@@
Marchesi et al. [2020] proposed a multi-arm bandit algorithm on top of the Gaussian processes and offers theoretical justification. Our work differentiates from two perspectives. First, Marchesi et al. [2020] focused on two-player zero-sum games, while our work allows multi-player normal-form games. Second, the regret analysis in Marchesi et al. [2020] relied on a suboptimal gap in the denominators of the regret bound. As discussed by Lattimore and Szepesvári [2020], the major problem with this dependency is that this gap, in practice, could be arbitrarily small and downgrade the practicality of the resulting regret analysis. At the same time, our theoretical results of the regret bound rely on the maximum mutual information of GP instead and are gap-independent.
@@ 2.2 Other online learning algorithms
Similar to the work by Aprem and Roberts [2021], Chapman et al. [2013] also studied convergence to Nash equilibria in potential games with rewards that are initially unknown. Different from the Bayesian optimization approach, they proposed a multi-agent version of Q-learning to estimate the reward functions using novel forms of the $\epsilon$–greedy learning policy. Jordan [1991] studied Bayesian learning of equilibrium, assuming each agent knows their utility information but not others. This work is also related to learning other equilibrium concepts in game theory and Bayesian optimization with multiple structured utility functions, we refer to Appendix A for more detailed discussions and comparisons.
@@ 4.3 Efficient High-dimensional Optimization through ROI Reduction
when setting ${\hat{\mathcal{X}}}^{0}=\mathcal{X}$. This means that the ROI identification is actually a hierarchical filtering of the search space and is accelerated by its continuing shrinkage. There is no guarantee of the ROI shrinkage rate, potentially making its performance unstable in High-Dimensional BO (HDBO) tasks. There are several potential solutions. There are chances to incorporate existing orthogonal HDBO techniques, including sparse GP [McIntire et al., 2016, Moss et al., 2023] and dimension reduction for BO [Song et al., 2022, Wang et al., 2016, Letham et al., 2020, Munteanu et al., 2019, Papenmeier et al., 2022]. However, the methods require additional structural assumptions that do not necessarily hold in NE discovery and, therefore, require cautiousness depending on the application.
@@ A.1 Learning Stackelberg games
More broadly, our research is also related to learning the equilibrium in the game theory. Besides the Nash equilibrium studied in this paper, another well-studied game is the Stackelberg game [Von Stengel and Zamir, 2004, Gan et al., 2023]. Specifically, Stackelberg games model a two-step sequential decision-making process between two agents, a leader and a follower. This canonical model for strategic leader-follower interactions has been adopted for many applications in the real world, such as contract design, optimal pricing, security resource allocation, and optimal traffic routing [Bolton and Dewatripont, 2004, Dawkins et al., 2021, Paruchuri et al., 2008, Roth et al., 2016, Roughgarden, 2001, Yang et al., 2014]. Learning the Stackelberg equilibrium has also been extensively studied in the literature [Letchford et al., 2009, Peng et al., 2019, Dawkins et al., 2022, Han et al., 2024], it would be interesting to study the learning of Stackelberg game equilibria via Gaussian Processes.